OpenStudy (anonymous):

$300 at 5% annual interest for 7 years Find the simple interest earned in the account. A)$214.29 B)$11.67 C)$420.00 D)$105.00

OpenStudy (agl202):

Formula of simple interest is S.I=Principal value * (1 + (rate/100))-Principal value S.I=300$*(1+(0.05))-300$ =15$ As it is for 7 years, S.I=7*15$ =$105

OpenStudy (anonymous):

Thanks

OpenStudy (agl202):

Medal? Ur welcome! :)

OpenStudy (anonymous):

and another formula is P*R*Tin which P=300,R=5%,T=7. so the answer is 105